### v2

# Sampling reduced density matrix to extract fine levels of entanglement spectrum

Abstract
Low-lying entanglement spectrum provides the quintessential fingerprint to identify the highly entangled quantum matter with topological and conformal field-theoretical properties. However, when the entangling region acquires long boundary with the environment, such as that between long coupled chains or in two or higher dimensions, there unfortunately exists no universal yet practical method to compute the entanglement spectra with affordable computational cost. Here we propose a new scheme to overcome such difficulty and successfully extract the low-lying fine entanglement spectrum (ES). We trace out the environment via quantum Monte Carlo simulation and diagonalize the reduced density matrix to gain the ES. We demonstrate the strength and reliability of our method through long coupled spin chains and answer its previous controversy. In addition, 2D examples have also been displayed which reveal the continuous symmetry breaking phase through a tower of states ES. Our simulation results, with unprecedentedly large system sizes, establish the practical computation scheme of the entanglement spectrum with a huge freedom degree of environment.
Introduction.- Quantum informatics and condensed matter physics have been increasingly cross-fertilizing each other in recent years Amico et al. (2008); Laflorencie (2016a). Within this trend, quantum entanglement was proposed to detect the field-theoretical and topological properties of quantum many-body systems Vidal et al. (2003); Korepin (2004); Kitaev and Preskill (2006); Levin and Wen (2006). For example, it usually offers the direct connection to the conformal field theory (CFT) and categorical description Calabrese and Lefevre (2008); Fradkin and Moore (2006); Nussinov and Ortiz (2009a, b); Casini and Huerta (2007); Ji and Wen (2019, 2020); Kong et al. (2020); Wu et al. (2021a); Zhao et al. (2021); Wu et al. (2021b). Beyond the entanglement entropy (EE), Li and Haldane proposed that the entanglement spectrum (ES) is a more fundamental entanglement characteristic to quantify the intrinsic information of many-body systems Li and Haldane (2008); Thomale et al. (2010); Poilblanc (2010); Schliemann and Läuchli (2012); Laflorencie (2016b). They also suggested a profound correspondence between the ES of an entangled region with the energy spectra on the related virtual edges, which is the so-called Li-Haldane conjecture. Thereafter, low-lying ES has been widely used as a fingerprint to investigate CFT and topology in highly entangled quantum matter Pollmann et al. (2010); Fidkowski (2010); Yao and Qi (2010); Qi et al. (2012); Canovi et al. (2014); Luitz et al. (2014a, b, c); Chung et al. (2014a); Pichler et al. (2016); Cirac et al. (2011a); Stojanović (2020); Guo (2021); Grover (2013); Assaad et al. (2014); Assaad (2015); Parisen Toldin and Assaad (2018); Yu et al. (2022); Roósz and Timm (2021); Roósz and Held (2022).
However, most of the ES studies so far have focused on (quasi) 1D systems numerically, due to the exponential growth of computation complexity and memory cost. The existing numerical methods such as the exact diagonalization (ED) and the density matrix renormalization group (DMRG) have significant limitations for approaching large entangling region. We note that the quantum Monte Carlo (QMC) is a powerful tool for studying large size and higher dimensional open quantum many-body systems Cai et al. (2014); Yan et al. (2018), which has been developed to extract the entanglement spectral function combined with numerical analytic continuation Yan and Meng (2023); Wu et al. (2023); Song et al. (2023); Liu et al. (2023); Assaad et al. (2014); Assaad (2015); Parisen Toldin and Assaad (2018); Shao and Sandvik (2023); Sandvik (1998); Beach (2004); Sandvik (2016); Syljuåsen (2008); Shao et al. (2017); Yan et al. (2021); Zhou et al. (2021). Though the method overcomes the exponential wall problem and can obtain large-scale entanglement spectral functions, it has rough precision and fails to distinguish the fine levels while the level structure always carries important information of CFT and topology. Moreover, the spectral function is also different from the full spectrum, which we will further compare to demonstrate that their characteristics sometimes may be totally different. The other way to extract ES through QMC is to reconstruct it according to the knowledge of all integer Rényi entropies Chung et al. (2014b) because in principle, all the integer Rényi entropies can be calculated by QMC. The difficulty of this scheme is that the calculation of $n$-th Rényi entropy itself is not an easy task especially for higher $n$ while the higher $n$-th entropy contains more information of low-lying levels.
Therefore, a practical scheme to extract fine low-lying ES with reduced computation complexity is urgent to be proposed. In this paper, we develop a protocol which can efficiently compute the ES via quantum Monte Carlo simulation for the entangling region with long boundaries and in higher dimensions. It can obtain fine low-lying levels of ES and is not limited to the freedom degree of the environment.
To demonstrate the strength of our method, a Heisenberg ladder in which we choose one chain as the entangling region has been investigated carefully. Compared with previous work studying the same system Poilblanc (2010); Läuchli and Schliemann (2012); Cirac et al. (2011b), we can not only address a much larger size with a groundbreaking record, but also clarify some misunderstandings in past work Yan and Meng (2023); Poilblanc (2010) through our large-scale calculations. Furthermore, we calculate a 2D Heisenberg model to reveal the nontrivial continuous symmetry breaking phase using the structure of tower of states in the ES. We have to emphasize that our method is not limited to these particular systems, yet it can be widely applied to any models which can be simulated via QMC.
Method.- In a quantum many-body system, the ES of a subsystem (entangling region) $A$ with the rest of the system $B$ (environment) is constructed via the reduced density matrix (RDM), defined as the partial trace of the total density matrix $\rho$ over a complete basis of $B$, i.e. $\rho_{A}=\mathrm{Tr}_{B}(\rho)$. The RDM $\rho_{A}$ can be interpreted as an effective thermodynamic density matrix $e^{-\mathcal{H}_{A}}$ through an entanglement Hamiltonian $\mathcal{H}_{A}$. The spectrum of the entanglement Hamiltonian is usually denoted as the ES.
Due to the exponential growth of the computation complexity and the memory cost, it is intractable for the existing numerical methods such as the exact diagonalization (ED) and the density matrix renormalization group (DMRG) to approach entangling regions with long boundaries and higher dimensions. Because of the finite computer memory, these numerical methods are limited to short boundary only. Another way is to extract the low-lying entanglement spectrum (ES) through the quantum Monte Carlo simulation combined with stochastic analytic continuation (SAC), both in bosonic Yan and Meng (2023); Liu et al. (2023); Li et al. (2023) and fermionic systems Assaad et al. (2014); Assaad (2015); Parisen Toldin and Assaad (2018). Especially in bosonic systems, the ES of spin ladders can be obtained even under a very long length $L=100$ Yan and Meng (2023). However, the ES obtained in this way can not show the fine levels but only the dispersion and weight of the spectral function. Nevertheless, the refined structure of ES, e.g., degeneracy, is important for identifying the CFT and topology.
The solution comes from QMC $+$ ED: Tracing the environment via QMC and obtaining the exact low-lying levels through ED. The weight of RDM element $|C_{A}\rangle\langle C^{\prime}_{A}|$ can be written in the path integral language as
${\rho_{A}}_{C_{A},C^{\prime}_{A}}=\sum_{\{C_{B}\}}\langle C_{A},C_{B}|e^{-\beta H}|C^{\prime}_{A},C_{B}\rangle,~{}{\beta\rightarrow\infty},$ (1)
where $C_{A},~{}C^{\prime}_{A}$ are the configurations of the RDM, and $C_{B}$ is the configurations of environment $B$.
It can be treated as a general partition function under a special boundary condition along the imaginary time axis (see Fig. 1(a)), which is much more convenient to be simulated by quantum Monte Carlo (QMC). We deal with this special partition function in Eq. (1) in the framework of stochastic series expansion (SSE) method Sandvik and Kurkijärvi (1991); Sandvik (1999); Syljuåsen and Sandvik (2002); Yan et al. (2019); Yan (2022). Of course it can be simulated by other path integral QMC Suzuki et al. (1977); Hirsch et al. (1982); Suzuki (1976); Blöte and Deng (2002); Huang et al. (2020); Fan et al. (2023). The only difference compared with the normal SSE is that we open the boundary of the imaginary time in the region $A$ and keep the periodic boundary condition to that of the environment $B$.
Refer to captionRefer to captionRefer to caption
Figure 1: (a) A path integral configuration of a reduced density matrix. The length of imaginary time is $\beta$, and the time boundary of $B$/$A$ is periodic/open; (b) Lattice of the Spin-1/2 Heisenberg ladder from an overhead view, with spins in $A$ and $B$ depicted by red and blue circles respectively; (c) Its phase diagram by denoting $J_{\text{leg}}=\cos\theta$ and $J_{\text{rung}}=\sin\theta$.
The value of each element in the RDM can be approached with the frequency of the samplings. It means that the ${\rho_{A}}_{C_{A},C^{\prime}_{A}}$ is proportional to $N_{C_{A},C^{\prime}_{A}}/N_{\text{total}}$, where the $N_{C_{A},C^{\prime}_{A}}$ is the amount of samplings that the upper/below imaginary time boundary configuration is $C_{A}/C^{\prime}_{A}$ and $N_{\text{total}}$ is the total amount of samplings  [^1] It’s worth noting that the sign should be considered if the weight is not positive..
Example 1: Antiferromagnetic (AFM) Spin-1/2 Heisenberg Ladder.- To demonstrate the power of our method, we compute the ES of two-leg Heisenberg ladder with $L=12,~{}16,~{}20,~{}24$ where the entangled boundary splits the ladder into two chains as shown in Fig. 1(b). The ES of this model has been carefully studied via ED for small sizes where $L=10,12,14$ in previous work Poilblanc (2010) and it has a well-known phase diagram (See Fig. 1(c)). The spins on the ladder are coupled through the nearest neighbor Heisenberg interactions, where $J_{\text{leg}}$ denotes the intra-chain strength along the leg and $J_{\text{rung}}$ denotes the inter-chain strength on the rung. We first simulate with $J_{\text{leg}}=1$ and $J_{\text{rung}}=1.732$ ($\theta=\pi/3$) at $\beta=100$. The reason for this choice of parameters is that $\beta=100$ ($T=J_{\text{leg}}/100$) is a temperature low enough for a gapped rung singlet (I) phase which has been studied with ED in Ref. Poilblanc (2010), therefore it is convenient for us to compare the results.
Figure 2: (a)The comparison of low lying entanglement excitation spectra obtained by pure ED and QMC+ED with different Monte Carlo steps. The red lines are the results obtained by pure ED and the black dots are the results obtained by QMC+ED. Here we choose $L=8$. (b) Entanglement excitation spectra versus total momenta $k$ in the chain direction for $\theta=\pi/3$.
Figure 3: (a) The fitting of velocity $v$ according to the numerical results with different size. (b)The extrapolation of $v$ with $1/L$. (c) Spectral function of the system with size $L=24$ and its fitting of $v$. (d) Entanglement spectrum versus total momenta $k$ in the chain direction for $\theta=2\pi/3$. The system size is $L=20$. (e) The entanglement spectral function of FM Heisenberg ladder with $L=20$. Its quadratic fitting curve near the gapless point. (f) The quadratic fitting curve in the full entanglement spectrum.
Firstly, we present a comparison of the spectrum obtained through QMC+ED and directly through ED with system size $L=8$, as illustrated in Fig. 2(a). It is evident that the spectrum obtained via QMC+ED approaches that of ED as the sample size increases, and higher energy levels require a greater number of sampling iterations. It is easy to be understood because the QMC of course samples the lower energy configurations with heavier weight preferentially, thus the low-lying levels converge first.
After demonstrating the power of this method, we want to answer some physical questions. For example, there is a contradiction for the ES in the rung singlet (I) phase of the Heisenberg ladder in previous works Poilblanc (2010); Yan and Meng (2023). The ES here is expected to bear the low-energy CFT structure, i.e., the ground level of ES, $\xi_{0}$, will scale as $\xi_{0}/L=e_{0}+d_{1}/L^{2}+\mathcal{O}(1/L^{3})$ where the $d_{1}=\pi cv/6$ according to the CFT predication with the central charge $c=1$ and $v$ is the velocity of ES near the gapless point Poilblanc (2010), i.e. the Cloiseaux-Pearson spectrum of the quantum spin chain, $v|\sin(k)|$ des Cloizeaux and Pearson (1962). The fitting for the fine levels of the ES by ED at $L=14$ (its largest size) gave $v\sim 2.36$ while the fitting for the rough spectral function of entanglement by QMC combined with SAC at $L=100$ pointed to a double value $v\sim 4.58$ Yan and Meng (2023). Therefore, the later one thought it is because the finite size effect of ES is serious. Since our method avoids the limitation of the environment size, we can calculate larger sizes to see how the $v$ scales with $L$. If the ES has obvious finite size effect, the $v$ certainly will increase following the size. In Fig. 2(b), we give the the low-lying levels of ES in different sizes which is much larger than the pure ED’s results, so that we can inspect the finite size effect here.
The spectra of the system with various system sizes obtained by QMC+ED is shown in Fig. 2(b), and the results are consistent with the ED results show in Ref. Poilblanc (2010). In the Fig. 3(a), we fit the velocities of different sizes via linear functions. All the lines are close to each other, representing that their velocities are similar. Furthermore, we draw a fitting line for the velocity $v$ of different size $L$ in Fig. 3(b), which scales to a value in the region $[2.40,2.45]$ at infinite size. It turns out that the velocity is almost unchanged with size $L$ and convergent to $\sim 2.41$, which is inconsistent to the result of spectral function by QMC+SAC where $v\sim 4.58$.
In order to show the spectral property of the system and avoid the difference between the ES and entanglement spectral function, we also calculate the spectral function of entanglement. For a physical observable denoted by $\mathcal{O}$, the spectral function $S(\omega)$ can be written as
$S(\omega)=\frac{1}{\pi}\sum_{m,n}e^{-\beta E_{n}}\left|\left\langle n\right|\mathcal{O}\left|m\right\rangle\right|^{2}\delta(\omega-\left[E_{m}-E_{n}\right]),$ (2)
where $\{|n\rangle\}$ and $\{E_{n}\}$ are the eigenstates and eigenvalues of the Hamiltonian $\mathcal{H}$ respectively. In the following, we choose $\mathcal{O}$ as the spin operator $S^{z}$, $|m\rangle$ as the ground state, and $E_{m}$ as the ground state energy. The choice is same as the Ref. Yan and Meng (2023) which is convenient for comparisons.
Fig. 3(c) shows the spectral function of the system with size $L=24$ and we also give the velocity $v\sim 2.29$. It reveals that the contradiction between the ES and entanglement spectral function is not because of the difference of the definition. We favor the reason that the settings of some constant are different between the two Refs. Poilblanc (2010); Yan and Meng (2023)
Example 2: Ferromagnetic (FM) Spin-1/2 Heisenberg Ladder.- Furthermore, we simulate the case with ferromagnetic $J_{\text{leg}}=-1$ and antiferromagnetic $J_{\text{rung}}=1.732$ at $\beta=100$ on the two-leg Heisenberg ladder for $L=8,12,16,20$ and compare the results with that in Ref. Poilblanc (2010), in which case the ladder is in the gapped rung singlet (II) phase.
According to Refs. H. (1931); Fogedby (1980); Haldane (1982); Yan and Meng (2023), the lowest-energy excitations at $k=2\pi/L$ corresponds to the one magnon state and behaves as $E=2J_{\text{eff}}{\sin}^{2}(k/2)$, where $J_{\text{eff}}$ is an effective chain coupling. The results in Ref. Poilblanc (2010) did not show the quadratic behavior due to the small system size. Other studies explained this non-quadratic dispersion via long-range boundary Hamiltonian Läuchli and Schliemann (2012); Cirac et al. (2011b); Laflorencie (2016b). Meanwhile, the QMC+SAC simlation at $L=100$ in the Ref. Yan and Meng (2023) shows the spectral function of entanglement is indeed a quadratic dispersion, which further demonstrated the finite size effect of ES here.
In order to check whether the loss of quadratic dispersion is because of the finite size effect, we further try to calculate the ES and its spectral function for larger sizes. Fig. 3(d) shows the results with $L=20$, in which there is no obvious quadratic dispersion near $k=0$. Actually, we haven’t seen an obvious change of the dispersion from the linear to the quadratic when increasing the size.
In fact, the answer of the contradiction comes from the difference between the ES and entanglement spectral function instead of the finite size effect. The results of the FM Heisenberg ladder indicates that the spectrum and spectral function exhibit distinct behaviors at low energy levels. As it is shown in Fig. 3(e), the ES in $L=20$ does not show a quadratic dispersion while the spectral function does. We try to draw the curve of the spectral function in the ES levels as shown in Fig. 3(f), which further demonstrates that the spectral function goes up quickly beyond the lowest levels of ES. Thus, the spectral function is not the descender line of full spectrum at all in this FM case which is opposite to the AFM case. This is actually the intrinsic reason that the FM ES has no quadratic dispersion. It also answers that why the top of the entanglement spectral function Yan and Meng (2023) is much larger than the top of the descender line of ES Poilblanc (2010).
Figure 4: (a) The AFM Heisenberg model on the square lattice. The dashed lines are to illustrate the bipartition into two subsystems. The yellow dashed lines illustrate the cutting method that $A$ is a ring and the red dashed lines illustrate the method that $A^{\prime}$ is a block. For each cutting method, the other part is denoted as $B$. We consider the square lattice model with size $20\times 20$. We show the entanglement spectrum of EH corresponding to (b) $A$ with length $L=20$ and (c) $A^{\prime}$ with size $4\times 4$. The TOS levels are connected by a red line. All the data are calculated in the total $S^{z}=0$ sector.
Two dimensional examples.- We then calculate a periodic boundary condition (PBC) $20\times 20$ square lattice AFM Heisenberg model with two different cut-geometries of $A$. If we choose the $A$ as a chain as the yellow dots ($A$ region) in Fig. 4(a) displays, it presents a linear dispersion ES reflecting the Néel order of this AFM model, as shown in Fig.4(b). What’s more, the linear low-lying levels is the famous tower of states (TOS) structure which reveals the continuous symmetry breaking here Kolley et al. (2013); Alba et al. (2013); Wietek et al. (2017); Laflorencie (2016b). Meanwhile, if we set $A$ as a $4\times 4$ region as the red dots ($A^{\prime}$ region) in Fig. 4(a), the ES also shows a TOS character as shown in Fig.4(c). Usually, the TOS holds in cornerless cutting cases, this is the first evidence supporting the TOS also arises in cornered cutting.
The advantage and disadvantage of this method.- We have to emphasize that the computational complexity of this method is exponential/polynomial for the freedom degree of $A$/$B$. Since we trace all the freedom degrees of $B$ and sample them, the complexity of $B$ part is as same as the normal QMC which increases in the power-law. At the same time, we sample all the freedom degree of $A$ without any trace operation, thus its complexity is proportional to the dimension of the RDM, which is exponentially increasing. Therefore, the advantage of this method is that it is basically not limited to the size of the environment and can obtain the RDM and fine levels of the ES.
However, because the important information is always contained in the low-lying ES, while the importance sampling of QMC just presents the low-lying levels first, actually we can set a cut-off for the sampling to avoid the additional cost for high levels. For example, if we only consider the first excited gap of the ES (the Schmidt gapBayat et al. (2014); Gray et al. (2018)), the sampling amount could be small. As shown in the Fig. 2(a), the first gap converges very quickly over the Monte Carlo sampling.
To compare with ED, this method can obtain $\rho_{A}$ in a much larger size. To compare with the QMC+SAC scheme, it is not limited to the cut-geometry between $A$ and $B$, and it can extract the fine levels of low-lying ES  [^2] In the QMC+SAC method, because the spectral function is gained from the imaginary time correlation function via numerical analytic continuation, it requires that $A$ should have periodic boundary condition to guarantee a well-defined momentum $k$. In other words, the analytic continuation depends closely on the cut form of the entanglement. However, our method can study any geometry for the $A$.. Similarly, although the Bisognano-Wichmann theorem gives the detailed form of entanglement Hamiltonian, it holds under some strict conditions (e.g., the entanglement region has no corner, the original Hamiltonian satisfies translation symmetry) and may lose the effectiveness in many lattice models Dalmonte et al. (2022); Mendes-Santos et al. (2019, 2020); Giudici et al. (2018).
Conclusions and outlooks.- We propose a practical scheme to extract the fine levels of entanglement spectrum from the method of quantum Monte Carlo simulation combined with exact diagonalization. It can extract the fine information of low-lying ES for large-scale quantum many-body systems whose environment has huge degrees of freedom. Using this method, we answer the previous contradiction for the entanglement spectrum in a Heisenberg ladder system. Furthermore, we calculate the ES of a 2D AFM Heisenberg model within two different cut-geometries. Both the cornered and cornerless cases show the TOS character which reflects the continuous symmetry breaking here.
In summary, the reduced density matrix can be obtained in larger size via QMC within this frame. Therefore, other observables such as von Neumann entropy and off-diagonal measurements can also be extracted in this way.
Acknowledgement.- We would like to thank Wei Zhu, Yin Tang, Xue-Feng Zhang, Shangqiang Ning and Han-Qing Wu for fruitful discussions. We acknowledge the start-up funding of Westlake University. The authors also acknowledge Beijng PARATERA Tech Co.,Ltd.(https://www.paratera.com/) for providing HPC resources that have contributed to the research results reported within this paper.